# The Fefferman metric for twistor CR manifolds and conformal geodesics in dimension three

### v3

@@ 4.6. Null geodesics of $g^{\mathrm{F}}$
$\dot{\gamma}^{2}$ $=i\eta^{j}(\dot{\zeta}_{j}-\Gamma_{kj}{}^{l}\zeta_{l}\dot{x}^{k})=ig(\eta,\nabla_{\dot{x}}\zeta),$
$\dot{\gamma}^{3}$ $=-\frac{i}{2}\Bigl{(}\overline{\zeta^{j}}(\dot{\zeta}_{j}-\Gamma_{kj}{}^{l}\zeta_{l}\dot{x}^{k})-\zeta^{j}(\overline{\dot{\zeta}_{j}}-\Gamma_{kj}{}^{l}\overline{\zeta_{l}}\dot{x}^{k})\Bigr{)}={\rm Im}\,g(\overline{\zeta},\nabla_{\dot{x}}\zeta)=-ig(\overline{\zeta},\nabla_{\dot{x}}\zeta),$
@@
(4.5) $g^{\rm F}(\dot{\gamma},\dot{\gamma})=2(\dot{\gamma}^{1}\dot{\gamma}^{\overline{2}}+\dot{\gamma}^{2}\dot{\gamma}^{\overline{1}})+2\dot{\gamma}^{0}\dot{\gamma}^{3}=0.$
@@
$\nabla^{\mathrm{F}}_{\dot{\gamma}}\dot{\gamma}^{2}$ $=\ddot{\gamma}^{2}+iP(\zeta,\eta)(\dot{\gamma}^{0})^{2}+i\bigl{(}P(\zeta,\overline{\zeta})-P(\eta,\eta)\bigr{)}\dot{\gamma}^{0}\dot{\gamma}^{1}+iP(\zeta,\zeta)\dot{\gamma}^{0}\dot{\gamma}^{\overline{1}}$
$\quad-iP(\overline{\zeta},\eta)(\dot{\gamma}^{1})^{2}-iP(\zeta,\eta)|\dot{\gamma}^{1}|^{2},$
$\nabla^{\mathrm{F}}_{\dot{\gamma}}\dot{\gamma}^{3}$ $=\ddot{\gamma}^{3}+iP(\overline{\zeta},\eta)\dot{\gamma}^{0}\dot{\gamma}^{1}-iP(\zeta,\eta)\dot{\gamma}^{0}\dot{\gamma}^{\overline{1}}+iP(\overline{\zeta},\overline{\zeta})(\dot{\gamma}^{1})^{2}-iP(\zeta,\zeta)(\dot{\gamma}^{\overline{1}})^{2}.$
@@
$\dot{\gamma}^{0}=\mathrm{const},\quad\dot{\gamma}^{1}=\mathrm{const},$
@@
(4.6) $\ddot{\gamma}^{2}+iP(\zeta,\eta)(\dot{\gamma}^{0})^{2}+i\bigl{(}P(\zeta,\overline{\zeta})-$ $P(\eta,\eta)\bigr{)}\dot{\gamma}^{0}\dot{\gamma}^{1}+iP(\zeta,\zeta)\dot{\gamma}^{0}\dot{\gamma}^{\overline{1}}$
$\quad-iP(\overline{\zeta},\eta)(\dot{\gamma}^{1})^{2}-iP(\zeta,\eta)|\dot{\gamma}^{1}|^{2}=0,$
(4.7) $\ddot{\gamma}^{3}+iP(\overline{\zeta},\eta)\dot{\gamma}^{0}\dot{\gamma}^{1}-iP(\zeta,\eta)\dot{\gamma}^{0}\dot{\gamma}^{\overline{1}}+iP(\overline{\zeta},\overline{\zeta})(\dot{\gamma}^{1})^{2}-iP(\zeta,\zeta)(\dot{\gamma}^{\overline{1}})^{2}=0.$
@@
$V$ $=g(V,\eta)\eta+g(V,\overline{\zeta})\zeta+g(V,\zeta)\overline{\zeta},$
$g(V,W)$ $=g(V,\eta)g(W,\eta)+g(V,\zeta)g(W,\overline{\zeta})+g(V,\overline{\zeta})g(W,\zeta)$
@@
(4.8) $\dot{x}=\dot{\gamma}^{0}\eta+\dot{\gamma}^{\overline{1}}\zeta+\dot{\gamma}^{1}\overline{\zeta},\quad|\dot{x}|^{2}=(\dot{\gamma}^{0})^{2}+2|\dot{\gamma}^{1}|^{2}.$
@@
(4.9) $g(\nabla_{\dot{x}}\nabla_{\dot{x}}\zeta,\overline{\zeta})$ $=i\ddot{\gamma}^{3}-|\dot{\gamma}^{2}|^{2}-(\dot{\gamma}^{3})^{2},$ $g(\nabla_{\dot{x}}\nabla_{\dot{x}}\zeta,\zeta)$ $=(\dot{\gamma}^{2})^{2},$
$g(\nabla_{\dot{x}}\nabla_{\dot{x}}\zeta,\eta)$ $=-i\ddot{\gamma}^{2}+\dot{\gamma}^{2}\dot{\gamma}^{3},$ $g(\nabla_{\dot{x}}\nabla_{\dot{x}}\eta,\zeta)$ $=i\ddot{\gamma}^{2}+\dot{\gamma}^{2}\dot{\gamma}^{3},$
$g(\nabla_{\dot{x}}\nabla_{\dot{x}}\eta,\eta)$ $=-2|\dot{\gamma}^{2}|^{2}.$
@@
$g(\nabla_{\dot{x}}\nabla_{\dot{x}}\eta,\zeta)$ $=\frac{d}{dt}g(\nabla_{\dot{x}}\eta,\zeta)-g(\nabla_{\dot{x}}\eta,\nabla_{\dot{x}}\zeta)=-\frac{d}{dt}g(\eta,\nabla_{\dot{x}}\zeta)-g(\nabla_{\dot{x}}\eta,\zeta)g(\nabla_{\dot{x}}\zeta,\overline{\zeta})$
$=-\frac{d}{dt}g(\eta,\nabla_{\dot{x}}\zeta)+g(\eta,\nabla_{\dot{x}}\zeta)g(\nabla_{\dot{x}}\zeta,\overline{\zeta})=i\ddot{\gamma}^{2}+\dot{\gamma}^{2}\dot{\gamma}^{3}.$
Now we are ready to prove
@@
(4.10) $\nabla_{\dot{x}}\nabla_{\dot{x}}\dot{x}-|\dot{x}|^{2}P(\dot{x})=-\bigl{(}2|\dot{\gamma}^{2}|^{2}+(\dot{\gamma}^{3})^{2}+P(\dot{x},\dot{x})\bigr{)}\dot{x}.$
@@
$\quad-\bigl{(}(\dot{\gamma}^{0})^{2}+2|\dot{\gamma}^{1}|^{2}\bigr{)}\bigl{(}\dot{\gamma}^{0}P(\eta)+\dot{\gamma}^{\overline{1}}P(\zeta)+\dot{\gamma}^{1}P(\overline{\zeta})\bigr{)}.$
@@
$g(\nabla_{\dot{x}}\nabla_{\dot{x}}\dot{x}-|\dot{x}|^{2}P(\dot{x}),\zeta)$ $=\dot{\gamma}^{0}g(\nabla_{\dot{x}}\nabla_{\dot{x}}\eta,\zeta)+\dot{\gamma}^{\overline{1}}g(\nabla_{\dot{x}}\nabla_{\dot{x}}\zeta,\zeta)+\dot{\gamma}^{1}g(\nabla_{\dot{x}}\nabla_{\dot{x}}\overline{\zeta},\zeta)$
$\quad-\bigl{(}(\dot{\gamma}^{0})^{2}+2|\dot{\gamma}^{1}|^{2}\bigr{)}\bigl{(}\dot{\gamma}^{0}P(\eta,\zeta)+\dot{\gamma}^{\overline{1}}P(\zeta,\zeta)+\dot{\gamma}^{1}P(\overline{\zeta},\zeta)\bigr{)}$
$=i\dot{\gamma}^{0}\ddot{\gamma}^{2}+\dot{\gamma}^{0}\dot{\gamma}^{2}\dot{\gamma}^{3}+\dot{\gamma}^{\overline{1}}(\dot{\gamma}^{2})^{2}-i\dot{\gamma}^{1}\ddot{\gamma}^{3}-\dot{\gamma}^{1}|\dot{\gamma}^{2}|^{2}-\dot{\gamma}^{1}(\dot{\gamma}^{3})^{2}$
$\quad-\bigl{(}(\dot{\gamma}^{0})^{2}+2|\dot{\gamma}^{1}|^{2}\bigr{)}\bigl{(}\dot{\gamma}^{0}P(\zeta,\eta)+\dot{\gamma}^{\overline{1}}P(\zeta,\zeta)+\dot{\gamma}^{1}P(\zeta,\overline{\zeta})\bigr{)}.$
@@
$\dot{\gamma}^{0}\dot{\gamma}^{2}\dot{\gamma}^{3}+\dot{\gamma}^{\overline{1}}(\dot{\gamma}^{2})^{2}-\dot{\gamma}^{1}|\dot{\gamma}^{2}|^{2}=-2\dot{\gamma}^{1}|\dot{\gamma}^{2}|^{2}.$
Moreover, the geodesic equations (4.6), (4.7) give
$i\dot{\gamma}^{0}\ddot{\gamma}^{2}$ $=P(\zeta,\eta)(\dot{\gamma}^{0})^{3}+\bigl{(}P(\zeta,\overline{\zeta})-P(\eta,\eta)\bigr{)}(\dot{\gamma}^{0})^{2}\dot{\gamma}^{1}+P(\zeta,\zeta)(\dot{\gamma}^{0})^{2}\dot{\gamma}^{\overline{1}}$
$\quad-P(\overline{\zeta},\eta)\dot{\gamma}^{0}(\dot{\gamma}^{1})^{2}-P(\zeta,\eta)\dot{\gamma}^{0}|\dot{\gamma}^{1}|^{2},$
$-i\dot{\gamma}^{1}\ddot{\gamma}^{3}$ $=-P(\overline{\zeta},\eta)\dot{\gamma}^{0}(\dot{\gamma}^{1})^{2}+P(\zeta,\eta)\dot{\gamma}^{0}|\dot{\gamma}^{1}|^{2}-P(\overline{\zeta},\overline{\zeta})(\dot{\gamma}^{1})^{3}+P(\zeta,\zeta)\dot{\gamma}^{\overline{1}}|\dot{\gamma}^{1}|^{2}.$
@@
$g(\nabla_{\dot{x}}\nabla_{\dot{x}}\dot{x}-|\dot{x}|^{2}P(\dot{x}),\zeta)$ $=-2\dot{\gamma}^{1}|\dot{\gamma}^{2}|^{2}-\dot{\gamma}^{1}(\dot{\gamma}^{3})^{2}-\dot{\gamma}^{1}P(\dot{x},\dot{x})$
$=-\bigl{(}2|\dot{\gamma}^{2}|^{2}+(\dot{\gamma}^{3})^{2}+P(\dot{x},\dot{x})\bigr{)}g(\dot{x},\zeta).$
@@
$g(\nabla_{\dot{x}}\nabla_{\dot{x}}\dot{x}-|\dot{x}|^{2}P(\dot{x}),\eta)$ $=\dot{\gamma}^{0}g(\nabla_{\dot{x}}\nabla_{\dot{x}}\eta,\eta)+\dot{\gamma}^{\overline{1}}g(\nabla_{\dot{x}}\nabla_{\dot{x}}\zeta,\eta)+\dot{\gamma}^{1}g(\nabla_{\dot{x}}\nabla_{\dot{x}}\overline{\zeta},\eta)$
$\quad-\bigl{(}(\dot{\gamma}^{0})^{2}+2|\dot{\gamma}^{1}|^{2}\bigr{)}\bigl{(}\dot{\gamma}^{0}P(\eta,\eta)+\dot{\gamma}^{\overline{1}}P(\zeta,\eta)+\dot{\gamma}^{1}P(\overline{\zeta},\eta)\bigr{)}$
$=-2\dot{\gamma}^{0}|\dot{\gamma}^{2}|^{2}-i\dot{\gamma}^{\overline{1}}\ddot{\gamma}^{2}+\dot{\gamma}^{\overline{1}}\dot{\gamma}^{2}\dot{\gamma}^{3}+i\dot{\gamma}^{1}\ddot{\gamma}^{\overline{2}}+\dot{\gamma}^{1}\dot{\gamma}^{\overline{2}}\dot{\gamma}^{3}$
$\quad-\bigl{(}(\dot{\gamma}^{0})^{2}+2|\dot{\gamma}^{1}|^{2}\bigr{)}\bigl{(}\dot{\gamma}^{0}P(\eta,\eta)+\dot{\gamma}^{\overline{1}}P(\zeta,\eta)+\dot{\gamma}^{1}P(\overline{\zeta},\eta)\bigr{)}.$
@@
$\dot{\gamma}^{1}\dot{\gamma}^{\overline{2}}\dot{\gamma}^{3}+\dot{\gamma}^{\overline{1}}\dot{\gamma}^{2}\dot{\gamma}^{3}=-\dot{\gamma}^{0}(\dot{\gamma}^{3})^{2}$
@@
$-i\dot{\gamma}^{\overline{1}}\ddot{\gamma}^{2}+i\dot{\gamma}^{1}\ddot{\gamma}^{\overline{2}}$ $=-P(\zeta,\eta)(\dot{\gamma}^{0})^{2}\dot{\gamma}^{\overline{1}}-P(\overline{\zeta},\eta)(\dot{\gamma}^{0})^{2}\dot{\gamma}^{1}-2\bigl{(}P(\zeta,\overline{\zeta})-P(\eta,\eta)\bigr{)}\dot{\gamma}^{0}|\dot{\gamma}^{1}|^{2}$
$\quad-P(\zeta,\zeta)\dot{\gamma}^{0}(\dot{\gamma}^{\overline{1}})^{2}-P(\overline{\zeta},\overline{\zeta})\dot{\gamma}^{0}(\dot{\gamma}^{1})^{2}+2P(\overline{\zeta},\eta)\dot{\gamma}^{1}|\dot{\gamma}^{1}|^{2}$
$\quad+2P(\zeta,\eta)\dot{\gamma}^{\overline{1}}|\dot{\gamma}^{1}|^{2}.$
@@
$g(\nabla_{\dot{x}}\nabla_{\dot{x}}\dot{x}-|\dot{x}|^{2}P(\dot{x}),\eta)$ $=-2\dot{\gamma}^{0}|\dot{\gamma}^{2}|^{2}-\dot{\gamma}^{0}(\dot{\gamma}^{3})^{2}-\dot{\gamma}^{0}P(\dot{x},\dot{x})$
$=-\bigl{(}2|\dot{\gamma}^{2}|^{2}+(\dot{\gamma}^{3})^{2}+P(\dot{x},\dot{x})\bigr{)}g(\dot{x},\eta).$
@@
As an application, we have the following
Theorem 4.11.
Let $M_{i}\ (i=1,2)$ be the twistor CR manifolds over conformal $3$-manifolds $(\Sigma_{i},[g_{i}])\ (i=1,2)$. If $\widehat{f}\colon M_{1}\to M_{2}$ is a bundle isomorphism which is a CR equivalent map, then the underlying map $f\colon(\Sigma_{1},[g_{1}])\to(\Sigma_{2},[g_{2}])$ is a conformal isomorphism.

### v4

@@ 4.6 Null geodesics of $\boldsymbol{g^{\mathrm{F}}}$
$\dot{\gamma}^{2}={\rm i}\eta^{j}\bigl(\dot{\zeta}_{j}-\Gamma_{kj}{}^{l}\zeta_{l}\dot{x}^{k}\bigr)={\rm i}g(\eta,\nabla_{\dot{x}}\zeta),$
$\dot{\gamma}^{3}=-\frac{{\rm i}}{2}\bigl(\overline{\zeta^{j}}\bigl(\dot{\zeta}_{j}-\Gamma_{kj}{}^{l}\zeta_{l}\dot{x}^{k}\bigr)-\zeta^{j}\bigl(\overline{\dot{\zeta}_{j}}-\Gamma_{kj}{}^{l}\overline{\zeta_{l}}\dot{x}^{k}\bigr)\bigr)=\operatorname{Im}g\bigl(\overline{\zeta},\nabla_{\dot{x}}\zeta\bigr)=-{\rm i}g\bigl(\overline{\zeta},\nabla_{\dot{x}}\zeta\bigr),$
@@
$g^{\rm F}(\dot{\gamma},\dot{\gamma})=2\bigl(\dot{\gamma}^{1}\dot{\gamma}^{\overline{2}}+\dot{\gamma}^{2}\dot{\gamma}^{\overline{1}}\bigr)+2\dot{\gamma}^{0}\dot{\gamma}^{3}=0.$ (4.5)
@@
$\nabla^{\mathrm{F}}_{\dot{\gamma}}\dot{\gamma}^{2}=\ddot{\gamma}^{2}+{\rm i}P(\zeta,\eta)\bigl(\dot{\gamma}^{0}\bigr)^{2}+{\rm i}\bigl(P\bigl(\zeta,\overline{\zeta}\bigr)-P(\eta,\eta)\bigr)\dot{\gamma}^{0}\dot{\gamma}^{1}+{\rm i}P(\zeta,\zeta)\dot{\gamma}^{0}\dot{\gamma}^{\overline{1}}$
[…]
$\nabla^{\mathrm{F}}_{\dot{\gamma}}\dot{\gamma}^{3}=\ddot{\gamma}^{3}+{\rm i}P\bigl(\overline{\zeta},\eta\bigr)\dot{\gamma}^{0}\dot{\gamma}^{1}-{\rm i}P(\zeta,\eta)\dot{\gamma}^{0}\dot{\gamma}^{\overline{1}}+{\rm i}P\bigl(\overline{\zeta},\overline{\zeta}\bigr)\bigl(\dot{\gamma}^{1}\bigr)^{2}-{\rm i}P(\zeta,\zeta)\bigl(\dot{\gamma}^{\overline{1}}\bigr)^{2}.$
@@
$\dot{\gamma}^{0}=\mathrm{const},\qquad\dot{\gamma}^{1}=\mathrm{const},$
@@
$\ddot{\gamma}^{2}+{\rm i}P(\zeta,\eta)\bigl(\dot{\gamma}^{0}\bigr)^{2}+{\rm i}\bigl(P\bigl(\zeta,\overline{\zeta}\bigr)-P(\eta,\eta)\bigr)\dot{\gamma}^{0}\dot{\gamma}^{1}+{\rm i}P(\zeta,\zeta)\dot{\gamma}^{0}\dot{\gamma}^{\overline{1}}$
$\qquad{}-{\rm i}P\bigl(\overline{\zeta},\eta\bigr)\bigl(\dot{\gamma}^{1}\bigr)^{2}-{\rm i}P(\zeta,\eta)|\dot{\gamma}^{1}|^{2}=0,$ (4.6)
$\ddot{\gamma}^{3}+{\rm i}P\bigl(\overline{\zeta},\eta\bigr)\dot{\gamma}^{0}\dot{\gamma}^{1}-{\rm i}P(\zeta,\eta)\dot{\gamma}^{0}\dot{\gamma}^{\overline{1}}+{\rm i}P\bigl(\overline{\zeta},\overline{\zeta}\bigr)\bigl(\dot{\gamma}^{1}\bigr)^{2}-{\rm i}P(\zeta,\zeta)\bigl(\dot{\gamma}^{\overline{1}}\bigr)^{2}=0.$ (4.7)
@@
$V=g(V,\eta)\eta+g\bigl(V,\overline{\zeta}\bigr)\zeta+g(V,\zeta)\overline{\zeta},$
$g(V,W)=g(V,\eta)g(W,\eta)+g(V,\zeta)g\bigl(W,\overline{\zeta}\bigr)+g\bigl(V,\overline{\zeta}\bigr)g(W,\zeta)$
@@
$\dot{x}=\dot{\gamma}^{0}\eta+\dot{\gamma}^{\overline{1}}\zeta+\dot{\gamma}^{1}\overline{\zeta},\qquad|\dot{x}|^{2}=\bigl(\dot{\gamma}^{0}\bigr)^{2}+2\bigl|\dot{\gamma}^{1}\bigr|^{2}.$ (4.8)
@@
$g\bigl(\nabla_{\dot{x}}\nabla_{\dot{x}}\zeta,\overline{\zeta}\bigr)={\rm i}\ddot{\gamma}^{3}-\bigl|\dot{\gamma}^{2}\bigr|^{2}-\bigl(\dot{\gamma}^{3}\bigr)^{2},\qquad$ $g(\nabla_{\dot{x}}\nabla_{\dot{x}}\zeta,\zeta)=\bigl(\dot{\gamma}^{2}\bigr)^{2},$
$g(\nabla_{\dot{x}}\nabla_{\dot{x}}\zeta,\eta)=-{\rm i}\ddot{\gamma}^{2}+\dot{\gamma}^{2}\dot{\gamma}^{3},\qquad$ $g(\nabla_{\dot{x}}\nabla_{\dot{x}}\eta,\zeta)={\rm i}\ddot{\gamma}^{2}+\dot{\gamma}^{2}\dot{\gamma}^{3},$
$g(\nabla_{\dot{x}}\nabla_{\dot{x}}\eta,\eta)=-2\bigl|\dot{\gamma}^{2}\bigr|^{2}.$ (4.9)
@@
$g(\nabla_{\dot{x}}\nabla_{\dot{x}}\eta,\zeta)$ $=\frac{{\rm d}}{{\rm d}t}g(\nabla_{\dot{x}}\eta,\zeta)-g(\nabla_{\dot{x}}\eta,\nabla_{\dot{x}}\zeta)=-\frac{{\rm d}}{{\rm d}t}g(\eta,\nabla_{\dot{x}}\zeta)-g(\nabla_{\dot{x}}\eta,\zeta)g\bigl(\nabla_{\dot{x}}\zeta,\overline{\zeta}\bigr)$
$=-\frac{{\rm d}}{{\rm d}t}g(\eta,\nabla_{\dot{x}}\zeta)+g(\eta,\nabla_{\dot{x}}\zeta)g\bigl(\nabla_{\dot{x}}\zeta,\overline{\zeta}\bigr)={\rm i}\ddot{\gamma}^{2}+\dot{\gamma}^{2}\dot{\gamma}^{3}.$
Now we are ready to prove the following.
@@
$\nabla_{\dot{x}}\nabla_{\dot{x}}\dot{x}-|\dot{x}|^{2}P(\dot{x})=-\bigl(2\bigl|\dot{\gamma}^{2}\bigr|^{2}+\bigl(\dot{\gamma}^{3}\bigr)^{2}+P(\dot{x},\dot{x})\bigr)\dot{x}.$ (4.10)
@@
${}{-}\,\bigl(\bigl(\dot{\gamma}^{0}\bigr)^{2}+2\bigl|\dot{\gamma}^{1}\bigr|^{2}\bigr)\bigl(\dot{\gamma}^{0}P(\eta)+\dot{\gamma}^{\overline{1}}P(\zeta)+\dot{\gamma}^{1}P(\overline{\zeta})\bigr).$
@@
$g\bigl(\nabla_{\dot{x}}\nabla_{\dot{x}}\dot{x}-|\dot{x}|^{2}P(\dot{x}),\zeta\bigr)={}$ $\dot{\gamma}^{0}g(\nabla_{\dot{x}}\nabla_{\dot{x}}\eta,\zeta)+\dot{\gamma}^{\overline{1}}g(\nabla_{\dot{x}}\nabla_{\dot{x}}\zeta,\zeta)+\dot{\gamma}^{1}g\bigl(\nabla_{\dot{x}}\nabla_{\dot{x}}\overline{\zeta},\zeta\bigr)$
${}{-}\,\bigl(\bigl(\dot{\gamma}^{0}\bigr)^{2}+2\bigl|\dot{\gamma}^{1}\bigr|^{2}\bigr)\bigl(\dot{\gamma}^{0}P(\eta,\zeta)+\dot{\gamma}^{\overline{1}}P(\zeta,\zeta)+\dot{\gamma}^{1}P\bigl(\overline{\zeta},\zeta\bigr)\bigr)$
$={}$ ${\rm i}\dot{\gamma}^{0}\ddot{\gamma}^{2}+\dot{\gamma}^{0}\dot{\gamma}^{2}\dot{\gamma}^{3}+\dot{\gamma}^{\overline{1}}\bigl(\dot{\gamma}^{2}\bigr)^{2}-{\rm i}\dot{\gamma}^{1}\ddot{\gamma}^{3}-\dot{\gamma}^{1}\bigl|\dot{\gamma}^{2}\bigr|^{2}-\dot{\gamma}^{1}\bigl(\dot{\gamma}^{3}\bigr)^{2}$
${}{-}\bigl(\bigl(\dot{\gamma}^{0}\bigr)^{2}+2|\dot{\gamma}^{1}|^{2}\bigr)\bigl(\dot{\gamma}^{0}P(\zeta,\eta)+\dot{\gamma}^{\overline{1}}P(\zeta,\zeta)+\dot{\gamma}^{1}P\bigl(\zeta,\overline{\zeta}\bigr)\bigr).$
@@
$\dot{\gamma}^{0}\dot{\gamma}^{2}\dot{\gamma}^{3}+\dot{\gamma}^{\overline{1}}\bigl(\dot{\gamma}^{2}\bigr)^{2}-\dot{\gamma}^{1}\bigl|\dot{\gamma}^{2}\bigr|^{2}=-2\dot{\gamma}^{1}\bigl|\dot{\gamma}^{2}\bigr|^{2}.$
Moreover, the geodesic equations (4.6) and (4.7) give
${\rm i}\dot{\gamma}^{0}\ddot{\gamma}^{2}=P(\zeta,\eta)\bigl(\dot{\gamma}^{0}\bigr)^{3}+\bigl(P\bigl(\zeta,\overline{\zeta}\bigr)-P(\eta,\eta)\bigr)\bigl(\dot{\gamma}^{0}\bigr)^{2}\dot{\gamma}^{1}+P(\zeta,\zeta)\bigl(\dot{\gamma}^{0}\bigr)^{2}\dot{\gamma}^{\overline{1}}$
[…]
$-{\rm i}\dot{\gamma}^{1}\ddot{\gamma}^{3}=-P\bigl(\overline{\zeta},\eta\bigr)\dot{\gamma}^{0}\bigl(\dot{\gamma}^{1}\bigr)^{2}+P(\zeta,\eta)\dot{\gamma}^{0}\bigl|\dot{\gamma}^{1}\bigr|^{2}-P\bigl(\overline{\zeta},\overline{\zeta}\bigr)\bigl(\dot{\gamma}^{1}\bigr)^{3}+P(\zeta,\zeta)\dot{\gamma}^{\overline{1}}\bigl|\dot{\gamma}^{1}\bigr|^{2}.$
@@
$g\bigl(\nabla_{\dot{x}}\nabla_{\dot{x}}\dot{x}-|\dot{x}|^{2}P(\dot{x}),\zeta\bigr)$ $=-2\dot{\gamma}^{1}\bigl|\dot{\gamma}^{2}\bigr|^{2}-\dot{\gamma}^{1}\bigl(\dot{\gamma}^{3}\bigr)^{2}-\dot{\gamma}^{1}P(\dot{x},\dot{x})$
$=-\bigl(2\bigl|\dot{\gamma}^{2}\bigr|^{2}+\bigl(\dot{\gamma}^{3}\bigr)^{2}+P(\dot{x},\dot{x})\bigr)g(\dot{x},\zeta).$
@@
$g\bigl(\nabla_{\dot{x}}\nabla_{\dot{x}}\dot{x}-|\dot{x}|^{2}P(\dot{x}),\eta\bigr)={}$ $\dot{\gamma}^{0}g(\nabla_{\dot{x}}\nabla_{\dot{x}}\eta,\eta)+\dot{\gamma}^{\overline{1}}g(\nabla_{\dot{x}}\nabla_{\dot{x}}\zeta,\eta)+\dot{\gamma}^{1}g\bigl(\nabla_{\dot{x}}\nabla_{\dot{x}}\overline{\zeta},\eta\bigr)$
${}{-}\bigl(\bigl(\dot{\gamma}^{0}\bigr)^{2}+2\bigl|\dot{\gamma}^{1}\bigr|^{2}\bigr)\bigl(\dot{\gamma}^{0}P(\eta,\eta)+\dot{\gamma}^{\overline{1}}P(\zeta,\eta)+\dot{\gamma}^{1}P\bigl(\overline{\zeta},\eta\bigr)\bigr)$
$={}$ ${-}2\dot{\gamma}^{0}\bigl|\dot{\gamma}^{2}\bigr|^{2}-{\rm i}\dot{\gamma}^{\overline{1}}\ddot{\gamma}^{2}+\dot{\gamma}^{\overline{1}}\dot{\gamma}^{2}\dot{\gamma}^{3}+{\rm i}\dot{\gamma}^{1}\ddot{\gamma}^{\overline{2}}+\dot{\gamma}^{1}\dot{\gamma}^{\overline{2}}\dot{\gamma}^{3}$
${}{-}\bigl(\bigl(\dot{\gamma}^{0}\bigr)^{2}+2\bigl|\dot{\gamma}^{1}\bigr|^{2}\bigr)\bigl(\dot{\gamma}^{0}P(\eta,\eta)+\dot{\gamma}^{\overline{1}}P(\zeta,\eta)+\dot{\gamma}^{1}P\bigl(\overline{\zeta},\eta\bigr)\bigr).$
@@
$\dot{\gamma}^{1}\dot{\gamma}^{\overline{2}}\dot{\gamma}^{3}+\dot{\gamma}^{\overline{1}}\dot{\gamma}^{2}\dot{\gamma}^{3}=-\dot{\gamma}^{0}\bigl(\dot{\gamma}^{3}\bigr)^{2}$
@@
$-{\rm i}\dot{\gamma}^{\overline{1}}\ddot{\gamma}^{2}+{\rm i}\dot{\gamma}^{1}\ddot{\gamma}^{\overline{2}}={}$ ${-}P(\zeta,\eta)\bigl(\dot{\gamma}^{0}\bigr)^{2}\dot{\gamma}^{\overline{1}}-P\bigl(\overline{\zeta},\eta\bigr)\bigl(\dot{\gamma}^{0}\bigr)^{2}\dot{\gamma}^{1}-2\bigl(P\bigl(\zeta,\overline{\zeta}\bigr)-P(\eta,\eta)\bigr)\dot{\gamma}^{0}\bigl|\dot{\gamma}^{1}\bigr|^{2}$
${}{-}\,P(\zeta,\zeta)\dot{\gamma}^{0}\bigl(\dot{\gamma}^{\overline{1}}\bigr)^{2}\!-P\bigl(\overline{\zeta},\overline{\zeta}\bigr)\dot{\gamma}^{0}\bigl(\dot{\gamma}^{1}\bigr)^{2}+2P\bigl(\overline{\zeta},\eta\bigr)\dot{\gamma}^{1}\bigl|\dot{\gamma}^{1}\bigr|^{2}+2P(\zeta,\eta)\dot{\gamma}^{\overline{1}}\bigl|\dot{\gamma}^{1}\bigr|^{2}.$
@@
$g\bigl(\nabla_{\dot{x}}\nabla_{\dot{x}}\dot{x}-|\dot{x}|^{2}P(\dot{x}),\eta\bigr)$ $=-2\dot{\gamma}^{0}\bigl|\dot{\gamma}^{2}\bigr|^{2}-\dot{\gamma}^{0}\bigl(\dot{\gamma}^{3}\bigr)^{2}-\dot{\gamma}^{0}P(\dot{x},\dot{x})$
$=-\bigl(2\bigl|\dot{\gamma}^{2}\bigr|^{2}+\bigl(\dot{\gamma}^{3}\bigr)^{2}+P(\dot{x},\dot{x})\bigr)g(\dot{x},\eta).$
@@
As an application, we have the following.
Theorem 4.11 (cf. [25]).
Let $M_{i}$, $i=1,2$ be the twistor CR manifolds over conformal $3$-manifolds $(\Sigma_{i},[g_{i}])$, $i=1,2$. If $\widehat{f}\colon M_{1}\to M_{2}$ is a bundle isomorphism which is a CR equivalent map, then the underlying map $f\colon(\Sigma_{1},[g_{1}])\to(\Sigma_{2},[g_{2}])$ is a conformal isomorphism.